# Investigating Imperceptibility of Adversarial Attacks on Tabular Data: An Empirical Analysis

### v1

@@ 4.4 Analysis of Imperceptibility using Qualitative Properties
In five evaluated datasets, only Compas dataset exhibits the apparent feature interdependency among features. Feature Age and Age Category capture similar information but serve different purposes. Age represents the exact age of an individual, while Age Category would group individuals into broader age ranges for analytical purposes. Ensuring consistency between the Age and Age Category features is essential for maintaining the integrity of the dataset. If the value of Age is modified, it logically follows that the corresponding Age Category should also be updated consistently to reflect this change. Table 9 illustrates that when changing Age from 80 or 83 to 18, Case #285 did not preserve the feature interdependency between Age and Age Category, as Age Category remained greater than 45. Conversely, in Case #501, the Age Category was perturbed to less than 25, corresponding to the modified value of Age. It suggests that the process of perturbation or modification applied to these features was not consistently handled across cases. Further, the consideration of feature interdependency is missing from the algorithm design of attacks.
5 Discussion and Limitations
5.1 A trade-off between imperceptibility and performance on adversarial attacks
We conduct experiments encompassing all four quantitative properties related to imperceptibility. All five adversarial attacks with different performance have similar sparsity levels, indicating that there is no correlation between attack success and sparsity levels. However, proximity, deviation, and sensitivity appear to display inconsistent behaviours in the different levels of attack success rate. To analyse the potential relationship between these three metrics and attack success rate, all adversarial example are divided into two groups: successful attack examples and unsuccessful attack examples. Box plots are then utilised to compare the distribution of each metric across these two groups of adversarial examples.
Figure 4: The box plot illustrates the relationship between proximity and attack success: 1) In bounded attacks, successful attacks tend to exhibit larger $\ell_{2}$ values compared to unsuccessful attacks; 2) For unbounded attacks, where proximity is constrained by the attack budget, the trade-off pattern is not evident; 3) The distributions of $\ell_{2}$ distance for some experiment settings appear extremely narrow, depicted as a thin bar in the plot.
Figure 5: The box plot illustrates the relationship between deviation and attack success: 1) In bounded attacks, successful attacks tend to exhibit higher deviation levels compared to unsuccessful attacks; 2) For unbounded attacks, the similar trade-off pattern is shown in some combination of models and datasets; 3) The distributions of deviation for some experiment settings appear extremely narrow, depicted as a thin bar in the plot.
Figure 6: The box plot illustrates the relationship between sensitivity and attack success: 1) In bounded attacks, successful attacks tend to exhibit higher sensitivity levels compared to unsuccessful attacks; 2) For unbounded attacks, the similar trade-off pattern is shown in some combination of models and datasets; 3) The distributions of sensitivity for some experiment settings appear extremely narrow, depicted as a thin bar in the plot.
As depicted in Figure 4, the distributions of the two groups of adversarial examples reveal that successful attack instances typically exhibit larger values in $\ell_{2}$ distance compared to unsuccessful attack examples, particularly evident in unbounded attacks. However, in bounded attacks, where there’s a constrained attack budget in place limiting the maximum proximity, the pattern is less distinct. The similar patterns are also identified in deviation (Figure 5) and sensitivity (Figure 6). The distributions of deviation and sensitivity for Deepfool, C&W attack and LowProFool shift from low to high when changing the attack success from false to true. For bounded attacks, although the limited attack budget still has an impact, trade-off patterns are observed in some experimental settings.
By incorporating the comparative analysis between attack success rate and three imperceptibility metrics, we identify there is a trade-off between effectiveness and imperceptibility for adversarial attacks on tabular data. It emphasises the intricate balance that must be maintained to develop both effective and imperceptible adversarial attacks.
5.2 No one silver bullet for attacking tabular data, but unbounded attacks demonstrate greater promise
@@
According to our evaluation on the effectiveness and imperceptibility of five adversarial attack algorithms, none of them can simultaneously achieve the highest attack success rate and the highest level of imperceptibility. Despite their high success rates, bounded attacks like FGSM and PGD often struggle to maintain high levels of imperceptibility, especially when considering metrics such as deviation and sensitivity. Additionally, determining an appropriate attack budget poses a significant challenge for bounded attacks on tabular data, as it directly impacts the trade-off between attack success and imperceptibility. In theory, by systematically investigating different budget values and evaluating their impact on both success rate and imperceptibility metrics, researchers can identify the sweet spot that maximises attack effectiveness while minimising perceptibility. However, conducting such an exhaustive search for the best attack budgets for each dataset is impractical due to resource constraints and the vast diversity of datasets.
@@
On the other hand, unbounded attacks demonstrate more promise in crafting imperceptible adversarial examples. C&W and DeepFool outperform FGSM and PGD in the evaluation of proximity, deviation and sensitivity. Nevertheless, further improvements are necessary for unbounded attacks to enhance their performance in attack success rate of attacking tabular dataset, especially for C&W attacks.
@@
Moreover, none of these algorithms take into account the requirements of feasibility, immutability, and feature interdependency in tabular data. Although auxiliary methods such as feature masks can be employed during the implementation of the attack algorithms, the design of these algorithms does not inherently consider these factors, making them theoretically unsuitable for crafting imperceptible adversarial examples for tabular data.
5.3 Considering sparsity in adversarial attack algorithm optimisation for tabular data
Our evaluation indicates that all attack methods target the alteration of all numerical features, akin to how image-based attacks perturb all pixels. But as discussed in Section 3, achieving imperceptibility in tabular data necessitates attacks to alter as few features as possible. Modifying fewer features is less likely to be noticeable to the human eye and can evade detection. Conversely, attacks that modify many features can significantly reduce the level of imperceptibility, thereby increasing the risk of detection. The $\ell_{0}$ norm is non-differentiable and, therefore, cannot serve as the sole optimisation metric (Croce and Hein, 2019). However, when incorporated sparsity as a part of attack algorithm optimisation, it could possibly govern the number of altered features.
5.4 Assumption on uniform feature importance
This research operates under the assumption that all features contribute equally to the predictive models, analogous to the notion that each pixel holds equal importance in images. However, in real-world tabular datasets, intricate inter-dependencies exist among various features. This observation suggests that the importance of different features is not uniform. Later in the research, there is a need for an extension into non-uniform adversarial attacks (Erdemir et al., 2021; Nandy et al., 2023). Addressing non-uniform adversarial challenges will contribute to a more comprehensive understanding of the robustness and generalisation capabilities of predictive models in practical applications.
5.5 Limitations
@@
Tabular data poses various challenges when encoding categorical variables, and the selection of an appropriate encoding method can have a substantial impact on model performance. The prevalent use of distance-based metrics in attack techniques for measuring perturbation size is a key consideration for selecting the one-hot encoding. While one-hot encoding is a common approach to handle categorical features, it concurrently amplifies the total number of features in a dataset. This expansion results from each unique category in a categorical feature being represented by a distinct binary feature, thereby introducing computational challenges, including the well-known curse of dimensionality (Borisov et al., 2021). This necessitates a careful consideration of how distance metrics are applied to both numerical and categorical features to ensure effective and meaningful perturbation measurements in the pursuit of imperceptibility. Exploring alternative distance metrics, such as Gower’s distance (Gower, 1971) and other categorical feature similarity measures (Cost and Salzberg, 1993; Le and Ho, 2005), holds promise for advancing the design of attack methods tailored to tabular data.
Limitations of parameter settings
The current experiment settings on evaluating the effectiveness and imperceptibility of adversarial attacks on tabular data faces several notable limitations. Primarily, the restricted size of the tabular datasets under investigation hinders the generalisability of findings to larger and more diverse datasets. Another limitation lies in the use of simplistic predictive models, predominantly simple neural networks, which may not fully capture the complexity of real-world tabular data scenarios. The research is further confined by a narrow spectrum of attack methods, which may omit vulnerabilities exploitable by alternative adversarial strategies. Additionally, the fixed attack settings, particularly the consistent setting of the bounded attack budget $\epsilon$ to 0.3, may not accurately reflect the dynamic and variable nature of adversarial scenarios, potentially overlooking nuanced insights into the model’s robustness. Varying the attack budget across the entire range from 0 to 1 could yield valuable insights into the imperceptibility of adversarial attacks on tabular data models. Overcoming these limitations is essential to enhance the reliability and applicability of the research findings in real-world settings.
@@ 6 Conclusion
In conclusion, this paper has made significant contributions in advancing our understanding of imperceptibility in adversarial machine learning on tabular data. The derivation of seven key imperceptibility properties, including proximity, sparsity, deviation, sensitivity, immutability, feasibility and feature interdependency, provides a comprehensive framework for characterising adversarial attacks in the context of tabular data. Furthermore, the practical insights derived from our empirical evaluation on these imperceptibility properties and metrics offer valuable guidance, such as the identified trade-offs between attack success and imperceptibility, the efficacy of optimisation-based attacks, and limitations in controlling sparsity, for developing robust and imperceptible adversarial attacks.
While our investigation of adversarial attacks on tabular data has provided valuable insights, several limitations and directions for future research should be acknowledged. The use of one-hot encoding increases dataset dimensionality, prompting a need for careful consideration of distance metrics for both numerical and categorical features. Future research should expand to larger datasets, sophisticated models, and a wider array of attack strategies. Additionally, the absence of the ablation study signals a need for future investigations into the impact of feature removal on adversarial robustness and imperceptibility.
@@
Zhipeng He: Conceptualization, Methodology, Software, Investigation, Writing – Original Draft, Visualization. Chun Ouyang: Conceptualization, Methodology, Investigation, Writing – Original Draft. Laith Alzubaid: Supervision. Alistair Barros: Supervision. Catarina Moreira: Supervision, Investigation, Writing – Original Draft.
@@ A.2 Feasible Feature Range for Feasibility
BMI Body Mss Index Less than 18.5 (Underweight)

### v3

@@ 4.4 Analysis of Imperceptibility using Qualitative Properties
Among the five mixed datasets, we consider the COMPAS dataset for analysis since the dataset exhibits apparent interdependencies among features. Specifically, the features Age and Age Category capture similar information but serve different purposes: Age represents an individual’s exact age, while Age Category groups individuals into broader age ranges for analytical purposes. Maintaining consistency between these two features is essential for preserving the dataset’s integrity. Therefore, if the value of Age is altered, the Age Category should be correspondingly updated to reflect this change accurately.
Revisiting the two adversarial examples of the DeepFool attack in Table 10, we observe the following issue in Case #285. When Age was changed from 80 to 18, the Age Category was not updated accordingly and remained at “Greater than 45”, breaking the interdependency between these two features. This discrepancy arises because DeepFool does not account for feature interdependencies when perturbing input data. Therefore, although in Case #501 the Age Category was adjusted consistently with the modified Age, this alignment is most likely coincidental.
5 Discussion
5.1 Imperceptibility of Bounded vs. Unbounded Attacks
@@ 5 Discussion
Despite their high attack success rates, the two bounded attacks—FGSM and PGD—often struggle to maintain good imperceptibility, especially when considering metrics such as deviation and sensitivity. While a pre-defined attack budget is a key constraint affecting attack performance, determining an appropriate budget for bounded attacks on tabular data presents a significant challenge. In theory, systematically exploring different attack budget values and evaluating their impact on both success rate and imperceptibility metrics could reveal the optimal balance that maximises attack effectiveness while minimising perceptibility. However, conducting an exhaustive search to determine the best attack budget for each dataset is impractical due to resource limitations and the vast diversity of datasets.
@@
In contrast, unbounded attacks are more likely to craft imperceptible adversarial examples. C&W and DeepFool outperform FGSM and PGD in the evaluation of proximity, deviation and sensitivity. Nevertheless, further improvements are necessary for the C&W attack to enhance their performance in attack success rate of attacking tabular dataset.
@@
Moreover, none of these algorithms systematically address the requirements of feasibility, immutability, and feature interdependency in tabular data. While the design of these algorithms does not inherently consider these factors, auxiliary methods, such as feature masks, can be used during the implementation of the attack algorithms to preserve immutable features or control features in feasible ranges. One potential improvement is to integrate these qualitative imperceptibility constraints directly into the algorithm’s design.
5.2 Adversarial Attacks’ Imperceptibility vs Effectiveness
According to our evaluation on the effectiveness and imperceptibility of five adversarial attack algorithms, none of them can achieve both the highest attack success rate and the best imperceptibility at the same time. In order to investigate the relationship between the imperceptibility and effectiveness of adversarial attacks on tabular data, we analyse the correlation between the attack success rate and four quantitative properties related to imperceptibility for all attacks. In general, all five adversarial attacks share similar sparsity values within the same dataset, attempting to perturb all numerical features in the dataset regardless of attacks’ effectiveness. For the other three imperceptibility properties—proximity, deviation and sensitivity—we categorise the adversarial examples into two groups: successful and unsuccessful examples. Box plots are then used to compare the distribution of individual metrics across the two groups of adversarial examples.
Figure 6: A series of box plots illustrating the distribution of proximity $\ell_{2}$ for both successful and unsuccessful attack examples. The box plots are horizontally aligned for the same dataset and vertically aligned for the same predictive model. In each box plot, the X-axis represents different adversarial attacks, while the Y-axis shows the distribution of proximity $\ell_{2}$. The blue boxes represent unsuccessful attacks, while the orange boxes indicate successful attacks. Due to insufficient effectiveness, the combination of C&W attack and LinearSVC is not considered and their results are marked with $\dagger$.
Figure 7: A series of box plots illustrating the distribution of deviation for both successful and unsuccessful attack examples. The box plots are horizontally aligned for the same dataset and vertically aligned for the same predictive model. In each box plot, the X-axis represents different adversarial attacks, while the Y-axis shows the distribution of deviation. The blue boxes represent unsuccessful attacks, while the orange boxes indicate successful attacks. Due to insufficient effectiveness, the combination of C&W attack and LinearSVC is not considered and their results are marked with $\dagger$.
Figure 8: A series of box plots illustrating the distribution of sensitivity for both successful and unsuccessful attack examples. The box plots are horizontally aligned for the same dataset and vertically aligned for the same predictive model. In each box plot, the X-axis represents different adversarial attacks, while the Y-axis shows the distribution of sensitivity. The blue boxes represent unsuccessful attacks, while the orange boxes indicate successful attacks. Due to insufficient effectiveness, the combination of C&W attack and LinearSVC is not considered and their results are marked with $\dagger$.
As depicted in Figure 6, the following insights regarding the relationship between proximity $\ell_{2}$ and attack success can be drawn:
  • 1.
    The high success rate on the Breast Cancer dataset leads to an imbalance between the number of successful and unsuccessful attack examples, making it unfair to compare their metric distributions. Therefore, this dataset is not consider in the following analysis and marked with $\ast$ in the visualisation.
  • 2.
    For the three unbounded attacks—DeepFool, C&W and LowProFool, successful attacks tend to exhibit larger proximity $\ell_{2}$ values compared to unsuccessful attacks.
  • 3.
    For two bounded attacks—FGSM and PGD, their proximity $\ell_{2}$ values are constrained by the attack budget capping the maximum proximity.
For proximity, the fixed attack budget in our experiments results in no difference in the proximity $\ell_{\infty}$ distribution between successful and unsuccessful examples for bounded attacks. The three unbounded attacks show similar results to those of proximity $\ell_{2}$, that is, successful attacks yield larger proximity $\ell_{\infty}$ values compared to unsuccessful attacks.
Similar patterns can be observed from the results of deviation (Figure 7) and sensitivity (Figure 8), respectively. For deviation, successful attacks tend to show larger values compared to unsuccessful attacks in the case of three unbounded attacks. This pattern also appears in two bounded attacks on the COMPAS and Diabetes datasets. Regarding sensitivity, the values for successful attacks are generally higher than those for unsuccessful attacks, across all datasets for unbounded attacks and on the COMPAS and Diabetes datasets for bounded attacks.
Based on the above comparative analysis between attack success rate and three imperceptibility metrics, there is a trade-off between the effectiveness and imperceptibility of adversarial attacks on tabular data.
5.3 Challenges and Limitations
@@
Due to the use of $\ell_{p}$-based distance metrics in attack algorithms, encoding categorical variables when attacking tabular data is challenging. For example, ordinal encoding (or label encoding) can introduce an unintended order among categorical features. This unintended order often has an impact on the prediction results, as the attack algorithms might incorrectly infer that the numeric distance between categories reflects a true difference in the data. Given the challenges associated with ordinal encoding, one-hot encoding is our preferred encoding method for encoding categorical variables. However, one-hot encoding can increase the total number of features in a dataset as each unique category in a categorical feature is represented by a separate binary feature. This can lead to a high-dimensional feature space, which can result in computational challenges such as the curse of dimensionality (Borisov et al., 2021). This necessitates a careful consideration of how distance metrics are applied to both numerical and categorical features to ensure effective and meaningful perturbation measurements in the pursuit of imperceptibility. Exploring alternative distance metrics, such as Gower’s distance (Gower, 1971) and other categorical feature similarity measures (Cost and Salzberg, 1993; Le and Ho, 2005), offers more options in encoding methods for categorical features in tabular data.
Focusing Solely on Proximity
Current adversarial attacks commonly focus on using the proximity of adversarial examples to the original inputs as the sole measure of imperceptibility. However, our evaluation highlights there are other essential properties for achieving imperceptible attacks but they are often overlooked by the evaluated attacks. Specifically, the design of current adversarial attacks does not adequately consider quantitative properties such as sparsity, sensitivity, and deviation, nor does it address qualitative aspects like immutability, feasibility, and feature interdependency. By integrating these properties into attack design, we can enhance the imperceptibility of adversarial attacks. For instance, our evaluation of sparsity indicates that all attack methods target the alteration of all numerical features, which are similar to how image-based attacks perturb all pixels. But as discussed in Section 3, achieving imperceptibility in tabular data necessitates attacks to alter as few features as possible. Modifying fewer features is less likely to be noticeable to the human eye and thus harder to detect. Conversely, attacks that modify many features can significantly reduce imperceptibility, thereby increasing the risk of detection.
Assumption on uniform feature importance
Most current adversarial attacks operate under the assumption that all features contribute equally to predictive models, similar to the idea that each pixel in an image holds equal importance. However, in tabular datasets, the importance of different features is often not uniform due to their heterogeneous nature. Later in the research, there is a need for an extension into non-uniform adversarial attacks (Erdemir et al., 2021; Nandy et al., 2023). Addressing non-uniform adversarial challenges will contribute to a more comprehensive understanding of the robustness and generalisation capabilities of predictive models in practical applications.
@@ 6 Conclusion
In this paper, we have proposed seven key properties for evaluating the imperceptibility of adversarial attacks on tabular data, which are proximity, sparsity, deviation, sensitivity, immutability, feasibility, and feature interdependency. These properties can be used to comprehensively characterise imperceptible adversarial attacks on tabular data. Furthermore, we have conducted an empirical evaluation of imperceptibility for five adversarial attacks using the proposed metrics. The evaluation reveals that, except for proximity, these attacks often overlook the proposed imperceptibility properties in their algorithm design, and there is a trade-off between the effectiveness and imperceptibility of adversarial attacks. The insights gained from this evaluation offer valuable guidance for developing effective and imperceptible adversarial attacks on tabular data.
Our investigation of adversarial attacks on tabular data has provided valuable insights informing clear directions for the next stage of research. The use of one-hot encoding increases dataset dimensionality, highlighting the need for careful consideration of distance metrics for both numerical and categorical features. Additionally, the non-uniform importance of features in tabular data should be more thoroughly investigated to understand how varying feature significance impacts the design and effectiveness of adversarial attacks. Future research should expand to larger datasets, more sophisticated models, and a wider range of attack strategies.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT for refining the language, enhancing clarity, and improving the overall coherence of the writing. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@
Zhipeng He: Conceptualization, Methodology, Software, Investigation, Writing – Original Draft, Visualization. Chun Ouyang: Conceptualization, Methodology, Investigation, Writing – Original Draft. Laith Alzubaidi: Supervision. Alistair Barros: Supervision. Catarina Moreira: Supervision, Investigation, Writing – Original Draft.
@@ A.2 Feasible Feature Range for Feasibility
BMI Body Mass Index Less than 18.5 (Underweight)